# Towards General Continuous Memory for Vision-Language Models

### v1

@@ Abstract
Language models (LMs) and their extension, vision-language models (VLMs), have achieved remarkable performance across various tasks. However, they still struggle with complex reasoning tasks that require multimodal or multilingual real world knowledge. To support such capabilities, an external memory system that can efficiently provide relevant multimodal information is essential. Existing approaches generally concatenate image and text tokens into a long sequence as memory, which, however, may drastically increase context length and even degrade performance. In contrast, we propose using continuous memory-a compact set of dense embeddings-to more effectively and efficiently represent multimodal and multilingual knowledge. Our key insight is that a VLM can serve as its own continuous memory encoder. We empirically show that this design improves performance on complex multimodal reasoning tasks. Building on this, we introduce a data-efficient and parameter-efficient method to fine-tune the VLM into a memory encoder, requiring only 1.2% of the model’s parameters and a small corpus of 15.6K self-synthesized samples. Our approach CoMEM utilizes VLM’s original capabilities to encode arbitrary multimodal and multilingual knowledge into just 8 continuous embeddings. Since the inference-time VLM remains frozen, our memory module is plug-and-play and can be flexibly integrated as needed. Extensive experiments across eight multimodal reasoning benchmarks demonstrate the effectiveness of our approach. Code and data is publicly released here https://github.com/WenyiWU0111/CoMEM/tree/main.
@@ 1 Introduction
Inspired by how humans offload facts, plans, and ideas to external repositories like notebooks or databases for on-demand access, it is promising to develop a general external memory [^1] In this paper, external memory denotes any detachable module or function that supplies the knowledge without changing LM parameters, in contrast to internal memory that embeds knowledge by modifying parameters. that contains useful world knowledge for augmenting VLMs (lewis2020rag, ; wang2023longmem, ). Early approaches directly concatenate the collected useful information into a long token sequence, and feeds it into VLMs (lewis2020rag, ; guu2020realm, ) e.g., retrieval-augmented generation (RAG) methods. However, multimodal representations demand significantly more input tokens (e.g., 8 to 11427 tokens per image in Qwen2.5-VL (bai2025qwen2.5-vl, )). Thus, simple concatenation would greatly increase the input length, making it difficult for the memory content to be used (yu2024visrag, ) (see the degradation in performance shown in Table. 1 after using RAG). To solve the token overload issue, token pruning methods have been proposed to remove unimportant in-context tokens (chen2024image, ; zhang2024sparsevlm, ). However, token pruning generally leads to incomplete contextual contents, which impedes the VLM’s ability to accurately understand and utilize the compressed information (wen2025token, ).
@@
Compared to discrete tokens, continuous embeddings naturally have stronger representation capability for complex data (elhage2021mathematical, ; olsson2022incontext, ; hao2024training, ). This advantage makes them particularly promising for memory encoder architectures designed to condense multimodal information into continuous representations. However, training such encoders faces two key challenges: (1) achieving generalizable compression ability across diverse multimodal inputs, and (2) maintaining semantic alignment with the VLM (jaegle2021perceiver, ). While large-scale training can improve performance, it greatly increases the training cost and becomes heavily sensitive to the training data distribution. For example, when dominated by simple cases or a single domain, the encoder tends to overfit and degrade generalization performance. (arpit2017closer, ) (grangier-iter-2022-trade, ).
@@
In this paper, we focus on efficiently training a general continuous memory encoder to effectively supply multimodal knowledge for VLMs. To avoid costly training for semantic alignment, it is essential to minimize the representation gap between the memory encoder and downstream VLMs before training. Therefore, a natural way is to use the VLM itself as the memory encoder. Our empirical study confirms that the VLM can serve as a memory encoder for itself without any additional training. Benefiting from the stacked self-attention mechanism, its generated continuous embeddings in each layer have already aggregated rich semantic information (vaswani2017attention, ) (clark2019bert, ). As shown in Fig. 1, a simple rule-based embedding selection strategy for constructing the memory can greatly boost the performance of VLMs in complex multimodal multilingual reasoning tasks.
@@ 2 Empirical Analysis with VLM as Memory Encoder
In this section, we conduct an empirical study to examine: (1) whether VLM can serve as a continuous memory encoder to compress multimodal information into compatible embeddings. (2) whether a few embeddings from the VLM can preserve key information to improve multimodal reasoning tasks.
@@ 2.1 Analysis Setup
To compare the effectiveness of different memory and context compression methods, we select three benchmarks: InfoSeek (chen2023can, ), OK-VQA (marino2019ok, ), and A-OKVQA (schwenk2022okvqa, ). These benchmarks contain complex visual questions that require both accurate visual entity identification and multi-step reasoning to derive the correct answer. Following existing work (caffagni2024wikillava, ) (qi2024rora, ), for each question, we utilize CLIP-based retriever (radford2021learning, ) to collect relevant top-10 multimodal knowledge items from a Wikipedia-based source dataset WiT (srinivasan2021wit, ) to construct the input data for the memory.
Table 1: Comparison of training-free memory methods. Bold indicates the best performance.
@@ 3.4 Discussion
In Table 2, we compare our method CoMEM with six closely related works: i.e., multimodal RAG (EchoSight echosight and ReflectiVA reflectiva ), context compression (xRAG (xrag, ) and VoCo-LLaMA vocolama ), and LLM memory methods (LONGMEM wang2023longmem and MA-LMM (he2024malmm, )). The comparison spans three dimensions: Properties, where we examine whether the method is continuous (Con.) and plug-and-play (PP.); Scenarios, evaluating support for multimodal (MM) and multilingual (ML) inputs; and Training Cost, which includes the amount of training data required and trainable parameters.
Table 2: Comparison of representative methods with our approach. "I", "T", and "V" denote Image, Text, and Video modalities, respectively.
Properties Scenarios Training Cost
Method Con. PP MM ML Data Params
EchoSight ✗T 900K 300M
ReflectiVA ✓I+T 6.82M 8B
xRAG ✗T 3M 40M
VoCo-LLaMA ✓I/V 665K 7B
LONGMEM ✗T 114M 558M
MA-LMM ✓V+T NA 200M
CoMEM ✓I+T 15K 200M
@@ 4.1 Experimental Setup
(1) InfoSeek and OVEN are constructed from Wikipedia and consist of challenging factual questions; (2) MRAG-Bench, OK-VQA, A-OKVQA, and ViQuAE focus on multimodal real-world, knowledge-intensive tasks; (3) CVQA and multilingual InfoSeek evaluate model’s ability to reason diverse linguistic and cultural contexts. Further details about benchmarks can be found in Appendix A.
@@
We compare our method against three types of baselines: (1) VLMs; (2) VLMs with vanilla RAG; (3) advanced RAG methods; covering a total of 18 different models.
@@ 6 Conclusion
In this paper, we first empirically demonstrated that a VLM can serve as the memory encoder for itself, which can convert the multimodal knowledge into continuous embeddings. Then, we proposed a data-efficient and parameter-efficient approach to fine-tune the VLM into a continuous memory encoder. Concretely, we only fine-tuned 1.2% of its parameters on only 15.6k self-synthesized samples, and then the memory can effectively encode any multimodal and multilingual knowledge into only 8 continuous embeddings. Since we do not modify the inference-time VLM, our memory can be flexibly integrated with the VLM when necessary. Extensive experiments on six English and two multilingual vision-reasoning benchmarks, demonstrate the superiority of our approach.
In the future, we will explore broader application of our approach on various complex reasoning and planning tasks. Besides, we will integrate our continuous memory mechanism into multimodal agents, and test its effectiveness in transmitting the knowledge among multiple LMs and VLMs.

### v2

@@ Abstract
Language models (LMs) and their extension, vision-language models (VLMs), have achieved remarkable performance across various tasks. However, they still struggle with complex reasoning tasks that require multimodal or multilingual real-world knowledge. To support such capabilities, an external memory system that can efficiently provide relevant multimodal information is essential. Existing approaches generally concatenate image and text tokens into a long sequence as memory, which, however, may drastically increase context length and even degrade performance. In contrast, we propose using continuous memory-a compact set of dense embeddings-to more effectively and efficiently represent multimodal and multilingual knowledge. Our key insight is that a VLM can serve as its own continuous memory encoder. We empirically show that this design improves performance on complex multimodal reasoning tasks. Building on this, we introduce a data-efficient and parameter-efficient method to fine-tune the VLM into a memory encoder, requiring only 1.2% of the model’s parameters and a small corpus of 15.6K self-synthesized samples. Our approach CoMEM utilizes VLM’s original capabilities to encode arbitrary multimodal and multilingual knowledge into just 8 continuous embeddings. Since the inference-time VLM remains frozen, our memory module is plug-and-play and can be flexibly integrated as needed. Extensive experiments across eight multimodal reasoning benchmarks demonstrate the effectiveness of our approach. Code and data is publicly released here https://github.com/WenyiWU0111/CoMEM/tree/main.
Figure 1: CoMEM architecture in comparison to the traditional RAG method.
@@ 1 Introduction
Inspired by how humans offload facts, plans, and ideas to external repositories like notebooks or databases for on-demand access, it is promising to develop a general external memory [^1] In this paper, external memory denotes any detachable module or function that supplies the knowledge without changing LM parameters, in contrast to internal memory that embeds knowledge by modifying parameters. that contains useful world knowledge for augmenting VLMs (lewis2020rag, ; wang2023longmem, ). Early approaches directly concatenate the collected useful information into a long token sequence, and feeds it into VLMs (lewis2020rag, ; guu2020realm, ) e.g., retrieval-augmented generation (RAG) methods. However, multimodal representations demand significantly more input tokens (e.g., 8 to 11427 tokens per image in Qwen2.5-VL (bai2025qwen2.5-vl, )). Thus, simple concatenation would greatly increase the input length, making it difficult for the memory content to be used (yu2024visrag, ) (see the degradation in performance shown in Table 2 after using RAG). To solve the token overload issue, token pruning methods have been proposed to remove unimportant in-context tokens (chen2024image, ; zhang2024sparsevlm, ). However, token pruning generally leads to incomplete contextual contents, which impedes the VLM’s ability to accurately understand and utilize the compressed information (wen2025token, ).
@@
Compared to discrete tokens, continuous embeddings naturally have stronger representation capability for complex data (elhage2021mathematical, ; olsson2022incontext, ; hao2024training, ). This advantage makes them particularly promising for memory encoder architectures designed to condense multimodal information into continuous representations. However, training such encoders faces two key challenges: (1) achieving generalizable compression ability across diverse multimodal inputs, and (2) maintaining semantic alignment with the VLM (jaegle2021perceiver, ). While large-scale training can improve performance, it greatly increases the training cost and becomes heavily sensitive to the training data distribution. For example, when dominated by simple cases or a single domain, the encoder tends to overfit and degrade generalization performance (arpit2017closer, ) (grangier-iter-2022-trade, ).
@@
In this paper, we focus on efficiently training a general continuous memory encoder to effectively supply multimodal knowledge for VLMs. To avoid costly training for semantic alignment, it is essential to minimize the representation gap between the memory encoder and downstream VLMs before training. Therefore, a natural way is to use the VLM itself as the memory encoder. Our empirical study confirms that the VLM can serve as a memory encoder for itself without any additional training. Benefiting from the stacked self-attention mechanism, its generated continuous embeddings in each layer have already aggregated rich semantic information (vaswani2017attention, ) (clark2019bert, ). As shown in Fig. 2, even a simple rule-based embedding selection strategy for constructing the memory can greatly boost the performance of VLMs in complex multimodal multilingual reasoning tasks, compared to RAGs.
@@
Table 1: Comparison between our approach and other representative line of work.
Properties Scenarios Training Cost
Category Method Continuous  Pulg-and-Play Multimodal Multilingual Data Parameters
Multimodal-RAG EchoSight echosight ✗Text 900K 300M
ReflectiVA reflectiva ✓Image+Text 6.82M 8B
RoRA-VLM roravlm ✓Image+Text 1M 7B
Context-Compression xRAG xrag ✗Text 3M 40M
KV-Distill Chari2025KVDistillNL ✗Text 500K+ 150M
VoCo-LLaMA vocolama ✓Image/Video 665K 7B
LM Memory LONGMEM wang2023longmem ✗Text 114M 558M
MA-LMM he2024malmm ✓Video+Text NA 200M
M+  wang2025mplus ✗Text 5M NA
MemGPT  memgpt2024packer ✗Text NA NA
CoMEM ✓Image+Text 15K 200M
@@ 2 Empirical Analysis with VLM as Memory Encoder
In this section, we conduct an empirical study to examine (1) whether VLM can serve as a continuous memory encoder to compress multimodal information into compatible embeddings, and (2) whether a few embeddings from the VLM can preserve key information to improve multimodal reasoning tasks.
@@ 2.1 Analysis Setup
To compare the effectiveness of different memory and context compression methods, we select three benchmarks: InfoSeek (chen2023can, ), OK-VQA (marino2019ok, ), and A-OKVQA (schwenk2022okvqa, ). These benchmarks contain complex visual questions that require both accurate visual entity identification and multi-step reasoning to derive the correct answer. Following existing work (caffagni2024wikillava, ) (roravlm, ), for each question, we utilize CLIP-based retriever (radford2021learning, ) to collect relevant top-10 multimodal knowledge items from a Wikipedia-based source dataset WiT (srinivasan2021wit, ) to construct the input data for the memory.
Table 2: Comparison of training-free memory methods. Bold indicates the best performance. For VLM-as-Memory here, we use the cache KV from a VLM without fine-tuning, which differs from the main method described in Section 3 and is intended for preliminary exploration.
@@ 3.4 Discussion
In Table 1, we compare our method CoMEM with ten closely related works: i.e., multimodal RAG (EchoSight echosight ReflectiVA reflectiva , and RoRA-VLM roravlm ), context compression (xRAG (xrag, ), KV-Distill (Chari2025KVDistillNL, ) and VoCo-LLaMA vocolama ), and LLM memory methods (LONGMEM wang2023longmem , MA-LMM (he2024malmm, ), M+ (wang2025mplus, ), and MemGPT (memgpt2024packer, )). The comparison spans three dimensions: Properties, where we examine whether the method is continuous and plug-and-play; Scenarios, evaluating support for multimodal and multilingual inputs; and Training Cost, which includes the amount of training data required and trainable parameters.
@@ 4.1 Experimental Setup
Note (1) InfoSeek and OVEN are constructed from Wikipedia and consist of challenging factual questions. (2) MRAG-Bench, OK-VQA, A-OKVQA, and ViQuAE focus on multimodal real-world, knowledge-intensive tasks. (3) CVQA and multilingual InfoSeek evaluate model’s ability to reason diverse linguistic and cultural contexts. Further details about benchmarks are in Appendix A.
@@
We compare our method against three types of baselines: (1) VLMs, (2) VLMs with vanilla RAG, and (3) advanced RAG methods, covering a total of 18 different models.
@@ 6 Conclusion
In this paper, we empirically demonstrate that a VLM can effectively serve as its own memory encoder, capable of converting multimodal knowledge into compact continuous embeddings. Building on this insight, we develop a data- and parameter-efficient method to fine-tune the VLM as a continuous memory encoder. Specifically, by updating only 1.2% of the model’s parameters using just 15.6k self-synthesized samples, the resulting memory module can encode diverse multimodal and multilingual knowledge into merely 8 continuous embeddings. Importantly, since the VLM remains unchanged during inference, our memory module can be seamlessly integrated or detached as needed. Extensive evaluations across six English and two multilingual vision-reasoning benchmarks demonstrate the effectiveness and versatility of our approach.
In future work, we plan to extend our approach to a wider range of complex reasoning and planning tasks. Additionally, we aim to integrate the continuous memory mechanism into multimodal agents and evaluate its effectiveness in facilitating knowledge transfer across multiple language and vision-language models.